How do you find the length when you know the width is 5 feet. The equation is 36=2w+2(3+2w)
Answer by fcabanski(1391)

Perimeter = 2W + 2L


The given equation shows L = 3+2W


If width = 5, then L = 3+2*5 = 3+10 = 13